OpenStudy (90890908967777878888):

20 f+10 A.f(2+10) B.2(f+5) C.2(5+10)

OpenStudy (alexandervonhumboldt2):

what do you need to do?

OpenStudy (90890908967777878888):

what

OpenStudy (alexandervonhumboldt2):

A.f(2+10)=2f+20 B.2(f+5)=2f+10 C.2(5+10)=10+20=30

OpenStudy (lochana):

can you provide more information about your question

OpenStudy (trantom):

or for C 2(5+10)=2(15)=30 if that's easier to think about

OpenStudy (90890908967777878888):

why

OpenStudy (alexandervonhumboldt2):

we cannot answer question if no enough info provided

OpenStudy (alexandervonhumboldt2):

it is impossible

OpenStudy (90890908967777878888):

yhalll kind confusing me a litttle

OpenStudy (alexandervonhumboldt2):

or you dont need any help and just spamming using fake question?

OpenStudy (trantom):

Your guys' names are so long xD

OpenStudy (90890908967777878888):

@AlexandervonHumboldt2 what u wwant me say about the thing

OpenStudy (trantom):

can you specify what "20 f+10" means?

OpenStudy (alexandervonhumboldt2):

@ganeshie8 @amistre64 @Zale101

OpenStudy (trantom):

wait what this question is closed how are we typing on it...

OpenStudy (alexandervonhumboldt2):

looks like he is spamming

OpenStudy (trantom):

oh ganeshie8!! I remember him he helped me out :)

OpenStudy (90890908967777878888):

go to my anthore i just made

OpenStudy (90890908967777878888):

@trantom go to y anthore quetsion